Two point charges q and 4q are held at separation r.At what distance the electric field due to them is zero

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
kq/x^2=k4q/(r-x)^2

r^2+x^2-2rx=4x^2
3x^2+2rx-r^2=0

x=-2r+sqrt(4r^2+12r^2)/6= -2r+4r/6=1/3r

sher mohammad
b.tech, iit delhi